# Knowledge Graph Guided Evaluation of Abstention Techniques

### v1

@@ E.3 Exploring Universality of Abstention for Concepts
We also explore whether pre-training data has an impact on the how abstention works for different concepts. For this, we examine correlations between abstention performances and frequency of the concept in a pre-training corpus. We determine frequencies from the Dolma corpus (Soldaini et al., 2024) using the WIMBD’s n-gram lookup tool (Elazar et al., 2024). We consider the aggregate frequencies of concept terms and its descendants as the frequency for the concept. Correlations between the performance across models and techniques with the frequencies of the concept in the pre-training data are visualized in Figure 8.

### v2

@@ C.1 Enforcing Abstention via Prompting
As highlighted in §3.1, we disambiguate context for the concept by including the list of ancestors.
@@ E.3 Exploring Universality of Abstention for Concepts
We also explore whether pre-training data has an impact on the how abstention works for different concepts. For this, we examine correlations between abstention performances and frequency of the concept in a pre-training corpus. We determine frequencies from the Dolma corpus (Soldaini and et al., 2024) using the WIMBD’s n-gram lookup tool (Elazar et al., 2024). We consider the aggregate frequencies of concept terms and its descendants as the frequency for the concept. Correlations between the performance across models and techniques with the frequencies of the concept in the pre-training data are visualized in Figure 8.
@@ E.4 Abstention Performance across Compositions of Concepts
E.5 Robustness of Abstention Techniques against Adversarial Perturbations
We conduct a preliminary experiment to study the robustness of abstention techniques when subject to adversarial perturbations. We use the top 10 performing jailbreaks from Wei et al. (2023) and report the worst case results. The results are given in Table 10. Due to computational constraints, we report results only for LLaMA-3.1 8B over a subset of 30 concepts.
Technique Abstention Rate General-ization Specificity
Prompt (ZS) $23.3$
(-$73.7$)
$17.8$
(-$52.7$)
$95.3$
(+$20.6$)
Prompt (FS+CoT) $67.0$
(-$30.0$)
$43.3$
(-$44.4$)
$89.0$
(+$22.0$)
Activation Steering $0.3$
(-$95.0$)
$0.0$
(-$65.6$)
$100$
(+$25.8$)
SFT $0.3$
(-$79.4$)
$0.0$
(-$67.8$)
$99.3$
(+$37.6$)
SFT+DPO $0.0$
(-$72.4$)
$0.0$ (
-$69.4$)
$99.3$ (+$31.4$)
Table 10: Evaluation of abstention techniques over LLaMa-3.1 when subjected to adversarial perturbations. Values in parentheses denote the percentage change in metrics from the initial results over this subset without perturbations.
We find that the abstention techniques we study are not adversarially robust, as we observe significant drops in abstention rates ($30$-$95$%) and generalization across all techniques. This also highlights the fragility of the models which lack guardrails to defend against such attacks.